The following information is provided about an open economy with a government. Use the information to answer the questions that follow:
C = 450 + 0.4Y I = 350
G = 150
X = 70
Z = 35 + 0.1Y T = 0.15Y
Yf = 1550
Q.2.1 Calculate the level of autonomous spending in this economy.
(2)
Q.2.2
Calculate the size of the multiplier
(Note: Round your answer to two decimal places)
(4)
Q.2.3
Calculate the equilibrium level of income (Hint: use the multiplier method)
(2)
© The Independent Institute of Education (Pty) Ltd 2020
Page 6 of 10

20
2020
Q.2.4
Q.2.5
Q.2.6
Question 3
Calculate the tax revenue to the government of this country when the economy (2) remains in equilibrium.
Calculate what the new equilibrium income should be if the government of this (6) country decides to cancel all taxes, implying the tax rate would now be 0%.
Before the government decreased the tax rate, how much of government spending was required to bring the economy to full employment?
1
Expert's answer
2020-12-07T07:15:30-0500

Q.2.1…Calculate the level of autonomous spending in this economy

Y= C+I+G+(X-Z)

Y= 450+0.4Y+350+150+70-35-0.1Y

Y=985+0.4(Y-T)-0.1Y

Y=985+0.4(Y-0.15Y)-0.1Y

Y=985+0.24Y

Autonomous spending= 985

Q.2.2

MULTIPLIER=

In the above, c is the marginal propensity to consume(mpc), m is the marginal propensity to import(mpm) and t is the tax rate

Multiplier=

Multiplier=

Multiplier=

Multiplier= 1.31

Q.2.3

Y=985+0.24Y

Y-0.24Y=985

=

Y=1296.0526

Equilibrium level of income=1296

Question 3

Tax revenue of the government when the economy remains in equilibrium

T=0.15Y

T=0.15(1296)

T=194.41

Equilibrium level when tax is reduced to 0

Tax is zero now and the other components are unchanged

New aggregate spending= C+I+G+NX

= 450+0.4(Y-0)+350+150+70-35-0.1Y

=985+0.3Y

At equilibrium, Y=AE(Aggregate Expenditure)

Y=985+0.3Y

Y-0.3Y= 985

=

Y=1407.14

How much of government spending was required to bring the economy to full employment

AE= 450+0.4(Y-0.15Y)+350+150+70-35-0.1Y

=985+0.4×0.85Y-0.1Y

= 985+0.34Y-0.1Y

=985+0.24Y

So at equilibrium

Y=985+0.24Y

Y-0.24Y=985

=

Y=1296.0526

Y**=1296.05

The gap between initital output and full employment output is Yf-Y**

1550-1296.05= 253.95.